OpenStudy (anonymous):

add and write in simplest form. 4y over y+2 + 8 over y+2

OpenStudy (angela210793):

when the denoms are the same all u have to do is operate with the numerator which means add them

OpenStudy (anonymous):

y^2 - 4y + 4 - 10 = 4 (y - 2)^2 = 14 y - 2 = +/- sqrt(14) y = 2 +/- sqrt(14)

OpenStudy (anonymous):

so add 4y and 8?

OpenStudy (anonymous):

\[\frac{ 4y+8 }{ y+2}?\]

OpenStudy (anonymous):

yes. that looks correct.

OpenStudy (anonymous):

Okay thanks :)

OpenStudy (anonymous):

you can simplify it further though.

OpenStudy (anonymous):

factorize 4y+8 (Find its GCF)

OpenStudy (anonymous):

\[\frac{ 2y+4 }{ y }?\]

OpenStudy (anonymous):

no.

OpenStudy (anonymous):

I am confused

OpenStudy (anonymous):

\[4(y+2) \over y+2\]

OpenStudy (anonymous):

ohhh so then you cancel y+2 and you end up with 4?

OpenStudy (anonymous):

yes. I believe so.

OpenStudy (anonymous):

Okay. thank you.

OpenStudy (anonymous):

you're welcome.